9. (5 Pts.) A star with a radius R, 109 m is rotating with initial angular velocity w 108 rad/s. The star collapses in on itself until it becomes a neutron star with the same mass, and a final radius Ry 10t m. Modeling the star as a uniform solid sphere, what is the final angular velocity of the star? (Hint: angular momentum is conserved)
Answer #1

Since angular momentum is conserverd :

Li = Lf

Ii\omegai = If\omegaf

((2/5) M Ri2\omegai ) = (2/5) M Rf2\omega f

Ri2 / Rf2​ = \omega f / \omega i

1018 / 108 =  \omegaf / 10-8

102 rad/s = \omega f